OpenStudy (anonymous):

Choose the point on the terminal side of theta. For theta= (7pi)/6

OpenStudy (anonymous):

solve the system: 2x-y+z=3, x+2y+z=12, 4x-3y+z=1

OpenStudy (anonymous):

excuse me...Please dont post other questions of unanswred questions...

OpenStudy (anonymous):

sorry first time using this! not sure how to use it!

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Its alright. Thats why I was being cordial :) no hard feelings

OpenStudy (anonymous):

So is anyone able to help me?

OpenStudy (anonymous):

:)

OpenStudy (jdoe0001):

got a Unit Circle? if not, just get one :)

OpenStudy (anonymous):

tell me how does this thing works pls.

OpenStudy (anonymous):

i do. but that didnt help

OpenStudy (anonymous):

you should have a box on the top left corner of your screen that says ask a question. ask it there...

OpenStudy (jdoe0001):

hmm, did you find \(\bf \cfrac{7\pi}{6}\) on the Unit Circle?

OpenStudy (anonymous):

yes. at 210 degreess

OpenStudy (anonymous):

k thanks

OpenStudy (jdoe0001):

any values there for cosine and sine or x and y?

OpenStudy (anonymous):

yes -sqrt3, and -1

OpenStudy (anonymous):

both over 2

OpenStudy (anonymous):

so would my answer be -sqrt3, -1?

OpenStudy (jdoe0001):

right, well, then that's the point, whose coordinate is \(\bf \left(-\cfrac{\sqrt{3}}{2},-\cfrac{1}{2}\right)\)

OpenStudy (anonymous):

:\\\what?

OpenStudy (jdoe0001):

 the point on the terminal side of \(\bf \cfrac{7\pi}{6}\)

OpenStudy (anonymous):

ok...so I had the right answer?

OpenStudy (anonymous):

because the (over two) portion isnt an option in the answers...

OpenStudy (jdoe0001):

form the Unit Circle, yes

OpenStudy (anonymous):

:D yay

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw